OpenStudy (anonymous):

Help me Evaluating please Evaluate 60 ÷ 12 + 2(13 - 7). a. 2.5 c. 22 b. 6 d. 17 I got 42... What am I doing wrong?

OpenStudy (michele_laino):

you have to compute the division first, namely, what is: 60:12=...?

OpenStudy (anonymous):

5?

OpenStudy (michele_laino):

ok! next, you have to compute this: 13-7=...?

OpenStudy (anonymous):

6

OpenStudy (michele_laino):

that's right, then you have to compute: 2*6=...?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

5,6,12

OpenStudy (michele_laino):

finally you have to compute: 5 + 12=...?

OpenStudy (anonymous):

17

OpenStudy (kash_thesmartguy):

\[60\div12+2(13-7)\]\[5+2(13-7)\]\[5+(26-14)\]\[Answer=17\] D.

OpenStudy (michele_laino):

that's right!

OpenStudy (pawanyadav):

Solve bracket first Then Divide and after that add

OpenStudy (anonymous):

Algebra is so confusing to me :/.

OpenStudy (pawanyadav):

So D is right

OpenStudy (anonymous):

Hello

OpenStudy (kash_thesmartguy):

It's easy.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What is easy?

OpenStudy (kash_thesmartguy):

If you want help, visit khanacademy.org

OpenStudy (michele_laino):

reassuming, here are our steps: 60 ÷ 12 + 2(13 - 7)= =5+2(13-7)= =5+ 2*6= =5+12= =17

OpenStudy (anonymous):

Kash, some people struggle with reading, I am good at reading, some people struggle with writting, I am good at writting, some people struggle in history, I am good in history, I can not figure out Algebra, I spent hours on hours trying

OpenStudy (anonymous):

Kash, I already was on that website all year.

OpenStudy (pawanyadav):

It's not late till now

OpenStudy (anonymous):

ive been in no website called like that :D

OpenStudy (anonymous):

like muffins

OpenStudy (anonymous):

i like muffins